By Changing the order of Integration and evaluate ∫0^{a}∫0^{√a^{2}-y^{2}}(x^{2}+y^{2})dxdy `
1
Expert's answer
2021-09-08T07:47:52-0400

"∫_0^{a}∫_0^{\sqrt{a^{2}-y^{2}}}(x^{2}+y^{2})dxdy \\
∫_{y=0}^{y=a}∫_{x=0}^{x=\sqrt{a^{2}-y^{2}}}(x^{2}+y^{2})dxdy \\"

From the above case, limits of integration are

"x= 0 \space to \space x= \sqrt{a^2-y^2}\\
y= 0 \space to \space y= a\\
 x= \sqrt{a^2-y^2} \implies y= \sqrt{a^2-x^2}\\"

So new limits of integration are

"∫_{x=0}^{x=a}∫_{y=0}^{y=\sqrt{a^{2}-x^{2}}}(x^{2}+y^{2})dydx \\
∫_{0}^{a}∫_{0}^{\sqrt{a^{2}-x^{2}}}(x^{2}+y^{2})dydx \\
=\int _0^{\sqrt{a^2-x^2}}x^2dy+\int _0^{\sqrt{a^2-x^2}}y^2dy\\
=\int _0^a\left(x^2\sqrt{a^2-x^2}+\frac{\left(a^2-x^2\right)^{\frac{3}{2}}}{3}\right)dx\\
=\int _0^ax^2\sqrt{a^2-x^2}dx+\int _0^a\frac{\left(a^2-x^2\right)^{\frac{3}{2}}}{3}dx\\
=\frac{\pi }{16}a^4+\frac{\pi a^4}{16}"